OpenStudy (anonymous):

A person is watching a boat from the top of a lighthouse. The boat is approaching the lighthouse directly. When first noticed the angle of depression to the boat is 18°33'. When the boat stops, the angle of depression is 51°33'. The lighthouse is 200 feet tall. How far did the boat travel from when it was first noticed until it stopped? Round your answer to the hundredths place.

OpenStudy (anas.p):

 you can find a by using\[\cos x = (200/a)\]Let (a-b) be equal to z and you can find (a-b) by using \[\cos y (200/z)\] now z-a will give you the distance moved.

OpenStudy (anonymous):

What is z?

OpenStudy (anas.p):

z=(a-b) i have mentioned it above.